HUBBLE SPACE TELESCOPE DAILY REPORT #2275 PERIOD COVERED: 0000Z (UTC) 12/21/98 - 0000Z (UTC) 12/22/98 Daily Status Report as of 356/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Six Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Eleven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. Near the beginning of the final iteration of this proposal, the status buffer messages mentioned in 2.3 and HSTAR 6824 occurred. Any possible effect on science is not currently known. Otherwise, the observations were completed as planned, and no anomalies were noted. 1.4 Completed WF/PC-2 6825 (Quantitative Measures of the Evolution of the Cluster Galaxy Population) The WF/PC-2 was used to make observations of galaxy MS0016+16A. The goal of this proposal is to understand the differential evolution of cluster and field galaxies over the redshift range 0.3 < Z < 0.6. A wave length calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Eight Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.7 Completed Four Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 7358 (The Unusual Absorption Line System of PG2302+029 -- Ejected or Intervening?) The Space Telescope Imaging Spectrograph (CCD, Ma1 and MA2) was used to make observations of the unusual absorption line systems PG2302+029 to help to identify the cause of this absorption system and allow us to determine whether the system is an extreme example of previously known classes of quasar absorption lines or represents an entirely new phenomenon. A wavelength calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed STIS/CCD/MA1/MA2 7305 (The Physical Characteristics of Small- Scale ISM Structure) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the star HD6457 in order to measure the physical characteristics of small-scale ISM structure. A wavelength calibration observation was also taken. The observations were completed as planned, and no anomalies were reported. 1.11 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.12 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.13 Completed STIS/CCD/MA2 7402 (Chromospheric Activity in Old Stars: Halo Subdwarfs' Variability) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the old sub-dwarf star HD184499 in order to study its chromospheric activity. All normal stars cool enough to have convection zones share a common minimum level of chromospheric activity which decreases with advancing spectral type; most show higher emission levels which vary with time. A wavelength calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Camera Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 6844 (M87 Globular Clusters {#6746 Parallel Field Follow-up }) The WF/PC-2 was used to make observations of the globular clusters which are located on the "south" end of galaxy M-87. Sub-arcsecond dithering was employed to reduce the effect of undersampling of the point spread function. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Two Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C67.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 29 Successful : 29 2.3 Operations Notes: The SSR EDAC error counter was cleared four times per ROP SR-1. Using ROP NS-5, engineering data report 0 was cleared. As described in HSTAR 6824, two NICMOS status buffer messages were received. At 355/223026Z, there was a TPG write error: the message read, "the error bit was set when sending TPG command data". Then, at 355/224730Z, there was a Detector 1 compare error: the message read, "timing pattern ID comparison failed to match". This is similar to the events of HSTAR 6814. The proposal presented in 1.3 may have been affected. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH